OpenStudy (anonymous):

Tutorial on Permutation and Combination. Part 2

OpenStudy (anonymous):

1. ‘r’ things can be selected out of ‘n’ identical things in only one way. 2. if \[\huge ^nP_r \] denotes the number of permutation of ‘n’ different things, taking ‘r’ at a time, then \[^{ n }P_{ r }=n(n-1)(n-2).......(n-r+1)=\frac { n! }{ (n-r)! } \] 3. Note that \[^{ n }P_{ n } = n! \] 4. If \[^nC_r\] denotes the number of combination of ‘n’ different things, taking ‘r’ @ a time then \[ ^{ n }C_{ r }=\frac { n! }{ r!(n-r)! } =\frac { ^{ n }P_{ r } }{ r! } \quad where\quad r\le \quad n;\quad n\quad \in\quad N\quad \& r\quad \in\quad W\] 5. The number of circular Permutation of ‘n’ different things taken all at a time is (n-1)! . If clockwise and anticlockwise circular permutation are considered to be same, then its \[\frac { (n-1)! }{ 2 } \] Remark : Number of circular permutation of ‘n’ things when ‘p’ alike and the rest different taken all at a time distinguishing clockwise and anticlockwise arrangement is \[\frac { (n-1)! }{ p! } \] 6. ‘n’ persons can sit around a circular table in (n-1)! Ways. 7. Given ‘n’ different objects, the number of ways of selecting at least one of them is, ^nC_1+^nC_2 +^nC_3+.......+^nC_N = 2^n-1. This can also be stated as the total number of combination of ‘n’ distinct things. 8. Total number of ways in which it is possible to make a selection by taking some or all out of ‘p+q+r+…….. things, where ‘p’ are alike one kind, ‘q’ alike of the second kind & so on is given by : ‘(p+1)(q+1)(r+!)………-1’ 9. Number of ways in which it is possible to make a selection of ‘m+p+n=N’ things, where ‘p’ are alike one kind ‘m’ alike the second kind & ‘n’ alike the third kind taken ‘r’ at a time given coefficient of x^r 10. Number of ways in which ‘n’ identical things may be distributed among ‘p’ persons if each person may receive none, one or more thing is; ^{ n+p-1 }C_n \[^{ n }C_{ r }=^{ n }C_{ n-1\quad };\quad ^nC_0 = \ ^nC_n=1 \] \[^{ n }C_{ x }=^{ n }C_{ y\quad }\rightarrow \quad x=y\quad or\quad x+y=\quad n\] \[^{ n }C_{ r }=^{ n }C_{ r-1\quad }=^{ n+1 }C_{ r } \] 11.\[ ^{ n }C_{ r } \] is maximum if (a) r=n/2, if n is even. (b) r=(n-1)/2 if n is odd

OpenStudy (anonymous):

thanks Lg

OpenStudy (anonymous):

@mathslover have a look

OpenStudy (anonymous):

@ParthKohli have a look

Parth (parthkohli):

Congratulations. http://www.math-for-all-grades.com/pandc.html

OpenStudy (anonymous):

then!

OpenStudy (anonymous):

here's my work

Parth (parthkohli):

'Your' work?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

Last 2 hrs I was working with the write and LaTeX

Parth (parthkohli):

http://www.math-for-all-grades.com/pandc.html See 'your' work here. Well, it's nice that you haven't *exactly* copied and pasted it.

OpenStudy (anonymous):

Man Every where the Concept is same

OpenStudy (anonymous):

right

Parth (parthkohli):

But the words are not exactly the same. Well, it's okay.

OpenStudy (anonymous):

If any doubt U can call my Sir . Dr. BK Parghania

OpenStudy (anonymous):

I go to his Coaching

OpenStudy (anonymous):

you took it bck

Parth (parthkohli):

I thought that you didn't want it bc you deleted that reply

OpenStudy (anonymous):

Sorry !!

OpenStudy (anonymous):

@ParthKohli @lgbasallote there's a user  when I click it I directly go to home! why ??